A basketball player has a history of making 80% of the foul shots taken during the games. What is the probability that one of the next five foul shots he
a. makes three missed shots?
b. makes at least 3 shots?
c. makes at most 2 shots?

Answer by stanbon(75887)

A basketball player has a history of making 80% of the foul shots taken during the games.
-----
Binomial Problem with p(make) = 0.8 ; and P(miss) = 0.2
--------------------------
What is the probability that of the next five foul shots he
a. misses 3 shots?
P(x = 3) = 5C3(0.2)^3*(0.8)^2 = binompdf(5,0.2,3) = 0.0512
------------------------------------
b. makes at least 3 shots?
P(3<= x <=5) = 1 - P(0<= x <=2) = 1-binomcdf(5,0.8,2) = 0.9421
-----
c. makes at most 2 shots?
P(0<= x<=2) = binomcdf(5,0.8,2) = 0.0579
=================
Cheers,
Stan H.
=================